# Long-Context Speculative Decoding with Efficient Drafting and Verification

### v1

@@ 3.3 Fast Tree Attention
Tree Speculative Decoding (Miao et al., 2024) leverages prefix trees and the causal structure of LLMs so that a draft model can propose multiple candidate sequences, while the target model only needs to verify them once, without altering the final results. In this process, Tree Attention plays a key role in ensuring both correctness and efficiency. Early works (Cai et al., 2024; Li et al., 2024) apply attention masks derived from prefix trees to the $QK^{\mathsf{T}}$ attention matrix, thus disabling wrong combinations between speculation tokens. However, these methods only run on PyTorch’s eager execution mode, precluding more advanced attention kernels (e.g., Flash_Decoding). As a result, the inference speed decreases significantly when the sequence length increases.
To address these performance bottlenecks, we propose a Hybrid Tree Attention mechanism, as illustrated in Figure 2. Our method is based on two key insights: 1) When performing Tree Attention, the queries and the cached key-value pairs $\{K_{\mathrm{cache}},V_{\mathrm{cache}}\}$ do not require additional masks; 2) Only the queries and the key-value pairs $\{K_{\mathrm{specs}},V_{\mathrm{specs}}\}$ from the current speculative tokens need masking, and the number of such speculative tokens is typically no more than 128. Based on these observations, we adopt a divide and aggregate approach that splits the attention computation into two parts and merges them afterward.
Splitting Key-Value Pairs. We partition all key-value pairs into two groups: $\{K_{\mathrm{cache}},V_{\mathrm{cache}}\}$: the cached part of the main sequence, which requires no attention mask; and $\{K_{\mathrm{specs}},V_{\mathrm{specs}}\}$: the speculative-stage part, which needs attention masks. For $\{K_{\mathrm{cache}},V_{\mathrm{cache}}\}$, we invoke the efficient Flash_Decoding kernel. For $\{K_{\mathrm{specs}},V_{\mathrm{specs}}\}$, we use our custom Triton kernel fused_mask_attn, which applies blockwise loading and masking in the KV dimension, enabling fast computation of attention. This step yields two sets of attention outputs $\{O_{\mathrm{cache}},O_{\mathrm{specs}}\}$ along with their corresponding denominators (i.e., log-sum-exp of all attention scores) $\{\mathrm{LSE}_{\mathrm{cache}},\mathrm{LSE}_{\mathrm{specs}}\}$.
@@
The theoretical guarantee is provided in Appendix A. As outlined above, this hybrid approach employs the highly efficient Flash_Decoding kernel for most of the computations in long-sequence inference and only uses a custom masking attention fused_mask_attn for the small number of speculative tokens. The kernel fused_mask_attn follows the design philosophy of Flash Attention 2 (Dao et al., 2023) by splitting $Q$, $K_{\text{specs}}$, and $V_{\text{specs}}$ into small blocks. This strategy reduces global memory I/O and fully leverages GPU streaming multiprocessors. Furthermore, for each block in the computation of $QK_{\text{specs}}^{\top}$, the mask matrix is loaded and used to apply the masking operation. The Hybrid Tree Attention effectively balances the parallel verification of multiple branches with improved inference speed, all without compromising the correctness.
@@ 3.4 Theoretical Analysis
Under regularity assumptions, the inference error between the Maximum Likelihood Estimate (MLE) trained by our method and the optimal GliDe parameter that take all tokens as inputs is upper bounded by
$\mathrm{A}\mathrm{E}+\mathrm{D}\mathrm{E}+\mathrm{G}\mathrm{E}$
with probability at least $1-\delta$. Here the approximation error is $\mathrm{A}\mathrm{E}=(1+d_{V}\exp(B))HB^{4}(1+B_{X}^{2}B^{2})\varepsilon$, the distribution shift error is $\mathrm{D}\mathrm{E}=\log(1+d_{V}\exp(B))\cdot\mathop{\mathrm{TV}}(\tilde{P}_{t},P_{t})$, and the generalization error is $\mathrm{G}\mathrm{E}=N^{-1/2}(d(d+d_{V})\log(1+NHB^{6})+\log\delta^{-1})$.
The formal statement and the proof are provided in Appendix D. The approximation error results from that we adopt a sliding window instead of using all tokens. This error is proportional to the attention score outside the window. The distribution shift error results from the positional embedding distributional shift. In our experiments, we set $\tilde{P}_{t}$ as the uniform distribution, which will have a smaller distribution shift than not adopting this position offset technique. The generalization error results from that we use $N$ samples to train the model.
Table 1: Mean accepted length ($\tau$), decoding speed (tokens/s), and speedups across different models and settings. Specifically, “Vanilla HF” refers to HuggingFace’s PyTorch-based attention implementation, while “Vanilla FA” employs Flash_Decoding. The speedup statistic calculates the acceleration ratio relative to the Vanilla HF method. All results are computed at $T=0$.
@@
Figure 3: Decoding speed (tokens/s) across different models and settings. All results are computed at $T=1$. The letters G, Q, M, L, and R on the horizontal axis represent the dataset GovReport, QMSum, Multi-News, LCC, and RepoBench-P respectively.
@@ 4.1 Settings
Target and draft models. We select four widely-used long-context LLMs, Vicuna (including 7B and 13B) (Chiang et al., 2023), LongChat (including 7B and 13B) (Li et al., 2023), LLaMA-3.1-8B-Instruct (Dubey et al., 2024), and QwQ-32B (Qwen, 2024), as target models. In order to make the draft model and target model more compatible, our draft model is consistent with the target model in various parameters such as the number of KV heads.
Training Process. We first train our draft model with Achor-Offest Indices on the SlimPajama-6B pretraining dataset (Soboleva et al., 2023). The random offset is set as a random integer from 0 to 15k for Vicuna models and LongChat-7B, and 0 to 30k for the other three models because they have longer maximum context length. Then we train our model on a small subset of the Prolong-64k long-context dataset (Gao et al., 2024) in order to gain the ability to handle long texts. Finally, we finetune our model on a self-built long-context supervised-finetuning (SFT) dataset to further improve the model performance. The position index of the last two stages is the vanilla indexing policy because the training data is sufficiently long. We apply flash noisy training during all three stages to mitigate the training and inference inconsistency, the extra overhead of flash noisy training is negligible. Standard cross-entropy is used to optimize the draft model while the parameters of the target model are kept frozen. To mitigate the VRAM peak caused by the computation of the logits, we use a fused-linear-and-cross-entropy loss implemented by the Liger Kernel (Hsu et al., 2024), which computes the LM head and the softmax function together and can greatly alleviate this problem. More details on model training can be found in Appendix B.
@@
We compare our method with the original target model and MagicDec, a simple prototype of TriForce. To highlight the significance of Flash_Decoding in long-context scenarios, we also present the performance of the original target model using both eager attention implemented by Huggingface and Flash_Decoding for comparison. To make a fair comparison, we also use Flash_Decoding for baseline MagicDec. The most important metric for speculative decoding is the walltime speedup ratio, which is the actual test speedup ratio relative to vanilla autoregressive decoding. We also test the average acceptance length $\tau$, i.e., the average number of tokens accepted per forward pass of the target LLM.
@@ 4.2 Main Results
Table 1 and Figure 3 show the decoding speeds and mean accept lengths across the five evaluated datasets at $T=0$ and $T=1$ respectively. Our proposed method significantly outperforms all other approaches on both summarization tasks and code completion tasks. When $T=0$, on summarization tasks, our method can achieve a mean accepted length of around 3.5 and a speedup of up to 2.67$\times$; and on code completion tasks, our method can achieve a mean accepted length of around 4 and a speedup of up to 3.26$\times$. This highlights the robustness and generalizability of our speculative decoding approach, particularly in long-text generation tasks. At $T=1$, our method’s performance achieves around 2.5$\times$ speedup, maintaining a substantial lead over MagicDec. This indicates that our approach is robust across different temperature settings, further validating its soundness and efficiency.
While MagicDec demonstrates competitive acceptance rates with LongSpec, its speedup is noticeably lower in our experiments. This is because MagicDec is primarily designed for scenarios with large batch sizes and tensor parallelism. In low-batch-size settings, its draft model leverages all parameters of the target model with sparse KV Cache becomes excessively heavy. This design choice leads to inefficiencies, as the draft model’s computational overhead outweighs its speculative benefits. Our results reveal that MagicDec only achieves acceleration ratios $>\!1$ on partial datasets when using a guess length $\gamma\!=\!2$ and consistently exhibits negative acceleration around 0.7$\times$ when $\gamma\!\geq\!3$, further underscoring the limitations of this method in such configurations.
Lastly, we can find attention implementation plays a critical role in long-context speculative decoding performance. In our experiments, “Vanilla HF” refers to HuggingFace’s attention implementation, while “Vanilla FA” employs Flash_Decoding. The latter demonstrates nearly a 2×2$\times$2 × speedup over the former, even as a standalone component, and our method can achieve up to 6×6$\times$6 × speedup over HF Attention on code completion datasets. This result underscores the necessity for speculative decoding methods to be compatible with optimized attention mechanisms like Flash_Decoding, especially in long-text settings. Our hybrid tree attention approach achieves this compatibility, allowing us to fully leverage the advantages of Flash_Decoding and further speedup.
4.3 Ablation Studies
Anchor-Offset Indices. The experimental results demonstrate the significant benefits of incorporating the Anchor-Offset Indices. Figure 4 shows that pretrained with Anchor-Offset Indices achieve a lower initial loss and final loss compared to those trained without it when training over the real long-context dataset. Notably, the initalization with Anchor-Offset Indices reaches the same loss level 3.93×3.93$\times$3.93 × faster than its counterpart. Table 2 further highlights the performance improvements across two datasets, a summary dataset Multi-News, and a code completion dataset RepoBench-P. Models with Anchor-Offset Indices exhibit faster output speed and larger average acceptance length $\tau$. These results underscore the effectiveness of Anchor-Offset Indices in enhancing both training efficiency and model performance.
@@ 4.3 Ablation Studies
Hybrid Tree Attention. The results presented in Figure 5 highlight the effectiveness of the proposed Hybrid Tree Attention, which combines Flash_Decoding with the Triton kernel fused_mask_attn. While the time spent on the draft model forward pass and the target model FFN computations remain comparable across the two methods, the hybrid approach exhibits a significant reduction in latency for the target model’s attention layer (the yellow part). Specifically, the attention computation latency decreases from 49.92 ms in the HF implementation to 12.54 ms in the hybrid approach, resulting in an approximately 75% improvement. The verification step time difference is minimal, further solidifying the conclusion that the primary performance gains stem from optimizing the attention mechanism.
@@
4.4 Long CoT Acceleration
Long Chain-of-Thought (LongCoT) tasks have gained significant attention recently due to their ability to enable models to perform complex reasoning and problem-solving over extended outputs (Qwen, 2024; OpenAI, 2024). In these tasks, while the prefix input is often relatively short, the generated output can be extremely long, posing unique challenges in terms of efficiency and token acceptance. Our method is particularly well-suited for addressing these challenges, effectively handling scenarios with long outputs. It is worth mentioning that MagicDec is not suitable for such long-output scenarios because the initial inference stage of the LongCoT task is not the same as the traditional long-context task. In LongCoT tasks, where the prefix is relatively short, the draft model in MagicDec will completely degrade into the target model, failing to achieve acceleration.
We evaluate our method on the QwQ-32B model using the widely-used benchmark AIME24 dataset, with a maximum output length set to 32k tokens. The results, illustrated in Figure 6, demonstrate a significant improvement in both generation speed and mean accepted tokens. Specifically, our method achieved a generation rate of 42.63 tokens/s, 2.25$\times$ higher than the baseline’s 18.92 tokens/s, and an average of 3.82 mean accepted tokens. Notably, QwQ-32B with LongSpec achieves even lower latency than the vanilla 7B model with Flash_Decoding, demonstrating that our method effectively accelerates the LongCoT model. These findings not only highlight the effectiveness of our method in the LongCoT task but also provide new insights into lossless inference acceleration for the o1-like model. We believe that speculative decoding will play a crucial role in accelerating this type of model in the future.
@@ 4.5 Throughput
The throughput results of Vicuna-7B on the RepoBench-P dataset show that LongSpec consistently outperforms both Vanilla and MagicDec across all batch sizes. At a batch size of 8, LongSpec achieves a throughput of 561.32 tokens/s, approximately 1.8$\times$ higher than MagicDec (310.58 tokens/s) and nearly 2$\times$ higher than Vanilla (286.96 tokens/s). MagicDec, designed with throughput optimization in mind, surpasses Vanilla as the batch size increases, reflecting its targeted improvements. However, LongSpec still sustains its advantage, maintaining superior throughput across all tested batch sizes.
@@ 5 Conclusion
In this paper, we propose LongSpec, a novel framework designed to enhance speculative decoding for long-context scenarios. Unlike previous speculative decoding methods that primarily focus on short-context settings, LongSpec directly addresses three key challenges: excessive memory overhead, inadequate training for large position indices, and inefficient tree attention computation. To mitigate memory constraints, we introduce an efficient draft model architecture that maintains a constant memory footprint by leveraging a combination of sliding window self-attention and cache-free cross-attention. To resolve the training limitations associated with short context data, we propose the Anchor-Offset Indices, ensuring that large positional indices are sufficiently trained even within short-sequence datasets. Finally, we introduce Hybrid Tree Attention, which efficiently integrates tree-based speculative decoding with Flash_Decoding. Extensive experiments demonstrate the effectiveness of LongSpec in long-context understanding tasks and real-world long reasoning tasks. Our findings highlight the importance of designing speculative decoding methods specifically tailored for long-context settings and pave the way for future research in efficient large-scale language model inference.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Correctness for Attention Aggregation

### v2

@@ 3.3 Fast Tree Attention
Tree Speculative Decoding (Miao et al., 2024) leverages speculation trees and the causal structure of LLMs so that a draft model can propose multiple candidate sequences, while the target model only needs to verify them once, without altering the final results. In this process, Tree Attention plays a key role in ensuring both correctness and efficiency. Early works (Cai et al., 2024; Li et al., 2024b) apply attention masks derived from prefix trees to the $QK^{\mathsf{T}}$ attention matrix, thus disabling wrong combinations between speculation tokens. However, these methods only run on PyTorch’s eager execution mode, precluding more advanced attention kernels such as Flash Attention. As a result, the inference speed decreases significantly when the sequence length increases.
To address these performance bottlenecks, we propose a Hybrid Tree Attention mechanism, as illustrated in Figure 2(c). Our method is based on two key insights: 1) when performing Tree Attention, as illustrated in the left part of Figure 2(c), the queries and the cached key-value pairs $\{K_{\mathrm{cache}},V_{\mathrm{cache}}\}$ do not require additional masks; 2) only the queries and the key-value pairs $\{K_{\mathrm{specs}},V_{\mathrm{specs}}\}$ from the current speculative tokens need masking as illustrated in the right part of Figure 2(c), and the number of such speculative tokens is typically small. Based on these observations, we adopt a divide and aggregate approach that splits the attention computation into two parts and merges them afterward.
Splitting Key-Value Pairs. We partition all key-value pairs into two groups: $\{K_{\mathrm{cache}},V_{\mathrm{cache}}\}$: the cached part of the main sequence, which requires no attention mask; and $\{K_{\mathrm{specs}},V_{\mathrm{specs}}\}$: the speculation-stage part, which needs attention masks. For $\{K_{\mathrm{cache}},V_{\mathrm{cache}}\}$, we invoke the efficient Flash Attention kernel. For $\{K_{\mathrm{specs}},V_{\mathrm{specs}}\}$, we use our custom Triton kernel fused_mask_attn, which applies blockwise loading and masking in the KV dimension, enabling fast computation of attention. This step yields two sets of attention outputs $\{O_{\mathrm{cache}},O_{\mathrm{specs}}\}$ along with their corresponding denominators (i.e., log-sum-exp of all attention scores) $\{\mathrm{LSE}_{\mathrm{cache}},\mathrm{LSE}_{\mathrm{specs}}\}$.
@@
The theoretical guarantee is provided in Appendix C. As outlined above, this hybrid approach employs the highly efficient Flash Attention kernel for most of the computations in long-sequence inference and only uses a custom masking attention fused_mask_attn for the small number of speculative tokens. The kernel fused_mask_attn follows the design philosophy of Flash Attention 2 (Dao et al., 2023) by splitting $Q$, $K_{\text{specs}}$, and $V_{\text{specs}}$ into small blocks. This strategy reduces global memory I/O and fully leverages GPU streaming multiprocessors. Furthermore, for each block in the computation of $QK_{\text{specs}}^{\top}$, the mask matrix is loaded and used to apply the masking operation. The Hybrid Tree Attention effectively balances the parallel verification of multiple branches with improved inference speed, all without compromising the correctness.
@@ 3.4 Theoretical Analysis
Under regularity assumptions, the inference error between the Maximum Likelihood Estimate (MLE) trained by our method and the optimal GliDe parameter that take all tokens as inputs is upper bounded by $\mathrm{A}\mathrm{E}+\mathrm{D}\mathrm{E}+\mathrm{G}\mathrm{E}$ with probability at least $1-\delta$. Here the approximation error is $\mathrm{A}\mathrm{E}=(1+d_{V}\exp(B))HB^{4}(1+B_{X}^{2}B^{2})\varepsilon$, the distribution shift error is $\mathrm{D}\mathrm{E}=\log(1+d_{V}\exp(B))\cdot\mathop{\mathrm{TV}}(\tilde{P}_{t},P_{t})$, and the generalization error is $\mathrm{G}\mathrm{E}=N^{-1/2}(d(d+d_{V})\log(1+NHB^{6})+\log\delta^{-1})$.
The formal statement and the proof are provided in Appendix D.2. The approximation error results from that we adopt a sliding window instead of using all tokens. This error is proportional to the attention score outside the window. The distribution shift error results from the positional embedding distributional shift. In our experiments, we set $\tilde{P}_{t}$ as the uniform distribution, which will have a smaller distribution shift than not adopting this position offset technique. The generalization error results from that we use $N$ samples to train the model.
Table 1: Average acceptance length $\tau$, decoding speed (tokens/s), and speedups across different models and settings. Specifically, “Vanilla HF” refers to HuggingFace’s PyTorch-based attention implementation, while “Vanilla FA” employs Flash Attention. The speedup statistic calculates the acceleration ratio relative to the Vanilla HF method. For the analysis of the reasons for the low speedup ratio of MagicDec, see Section 4.2 and 4.5. All results are computed at $T=0$.
@@
Figure 3: Decoding speed (tokens/s) across different models and settings. All results are computed at $T=1$. The letters G, Q, M, L, and R on the horizontal axis represent the datasets GovReport, QMSum, Multi-News, LCC, and RepoBench-P respectively.
@@ 4.1 Settings
Target and draft models. We select four widely-used long-context LLMs, Vicuna (including 7B and 13B) (Chiang et al., 2023), LongChat (including 7B and 13B) (Li et al., 2023), LLaMA-3.1-8B-Instruct (Dubey et al., 2024), and QwQ-32B (Qwen, 2024), as target models. In order to make the draft model and target model more compatible, our draft model is consistent with the target model in various parameters, such as the number of KV heads.
Training Process. We first train our draft model with Anchor-Offset Indices on the SlimPajama-6B pretraining dataset (Soboleva et al., 2023). The random offset is set as a random integer from 0 to 15k for Vicuna models and LongChat-7B, and 0 to 30k for the other three models because they have longer maximum context length. Then we train our model on a small subset of the Prolong-64k long-context dataset (Gao et al., 2024) in order to gain the ability to handle long texts. Finally, we finetune our model on a self-built long-context supervised-finetuning (SFT) dataset to further improve the model performance. The position index of the last two stages is the vanilla indexing policy because the training data is sufficiently long. We apply flash noisy training during all three stages to mitigate the training and inference inconsistency and the extra overhead of flash noisy training is negligible. More details on model training can be found in Appendix E.
@@
We compare our method with the original target model and MagicDec, a simple prototype of TriForce. To highlight the significance of Flash Attention in long-context scenarios, we also present the performance of the original target model using both eager attention implemented by Huggingface and Flash Attention for comparison. To make a fair comparison, we also use Flash Attention for baseline MagicDec. The most important metric for speculative decoding is the walltime speedup ratio, which is the actual test speedup ratio relative to vanilla autoregressive decoding. We also test the average acceptance length $\tau$, i.e., the average number of tokens accepted per forward pass of the target LLM.
@@ 4.2 Main Results
Table 1 and Figure 3 show the decoding speeds and average acceptance lengths across the five evaluated datasets at $T=0$ and $T=1$, respectively. Our proposed method significantly outperforms all other approaches on both summarization tasks and code completion tasks. When $T=0$, on summarization tasks, our method can achieve an average acceptance length of around 3.5 and a speedup of up to 2.67$\times$; and on code completion tasks, our method can achieve an average acceptance length of around 4 and a speedup of up to 3.26$\times$. This highlights the robustness and generalizability of our speculative decoding approach, particularly in long-text generation tasks. At $T=1$, our method’s performance achieves around 2.5$\times$ speedup, maintaining a substantial lead over MagicDec. This indicates that our approach is robust across different temperature settings, further validating its soundness and efficiency.
While MagicDec demonstrates competitive acceptance rates with LongSpec, its speedup is noticeably lower in our experiments. This is because MagicDec is primarily designed for scenarios with large batch sizes and tensor parallelism. In low-batch-size settings, its draft model leverages all parameters of the target model with sparse KV Cache becomes excessively heavy. This design choice leads to inefficiencies, as the draft model’s computational overhead outweighs its speculative benefits. Our results reveal that MagicDec only achieves acceleration ratios $>\!1$ on partial datasets when using a guess length $\gamma\!=\!2$ and consistently exhibits negative acceleration around 0.7$\times$ when $\gamma\!\geq\!3$, further underscoring the limitations of this method in such configurations. The performance of MagicDec in larger batch sizes can be found in Section 4.5.
@@
Lastly, we find that attention implementation plays a critical role in long-context speculative decoding performance. In our experiments, “Vanilla HF” refers to HuggingFace’s attention implementation, while “Vanilla FA” employs Flash Attention. The latter demonstrates nearly a 2×2$\times$2 × speedup over the former, even as a standalone component, and our method can achieve up to 6×6$\times$6 × speedup over HF Attention on code completion datasets. This result underscores the necessity for speculative decoding methods to be compatible with optimized attention mechanisms like Flash Attention, especially in long-text settings. Our hybrid tree attention approach achieves this compatibility, allowing us to fully leverage the advantages of Flash Attention and further speedup.
4.3 Ablation Studies
Anchor-Offset Indices. The experimental results demonstrate the significant benefits of incorporating the Anchor-Offset Indices. Figure 4 shows that the model trained with Anchor-Offset Indices achieves a lower initial loss and final loss compared to the one trained without them when training on the real long-context dataset. Notably, the initialization with Anchor-Offset Indices reaches the same loss level 3.93×3.93$\times$3.93 × faster than its counterpart. Table 2 further highlights the performance improvements across two datasets, a summary dataset Multi-News, and a code completion dataset RepoBench-P. Models with Anchor-Offset Indices exhibit faster output speed and larger average acceptance length $\tau$. These results underscore the effectiveness of Anchor-Offset Indices in enhancing both training efficiency and model performance.
Hybrid Tree Attention. The results presented in Figure 5 highlight the effectiveness of the proposed Hybrid Tree Attention, which combines Flash Attention with the Triton kernel fused_mask_attn. While the time spent on the draft model forward pass and the target model FFN computations remain comparable across the two methods, the hybrid approach exhibits a significant reduction in latency for the target model’s attention layer (the yellow part). Specifically, the attention computation latency decreases from 49.92 ms in the HF implementation to 12.54 ms in the hybrid approach, resulting in an approximately 75% improvement. The verification step time difference is minimal, further solidifying the conclusion that the primary performance gains stem from optimizing the attention mechanism.
@@ 4.3 Ablation Studies
4.4 Long Reasoning Acceleration
Long reasoning tasks have gained significant attention recently due to their ability to enable models to perform complex reasoning and problem-solving over extended outputs (Qwen, 2024; OpenAI, 2024). In these tasks, while the prefix input is often relatively short, the generated output can be extremely long, posing unique challenges in terms of efficiency and token acceptance. Our method is particularly well-suited for addressing these challenges, effectively handling scenarios with long outputs. It is worth mentioning that MagicDec is not suitable for such long-output scenarios because the initial inference stage of the long reasoning task is not the same as the traditional long-context task. In long reasoning tasks, where the prefix is relatively short, the draft model in MagicDec will completely degrade into the target model, failing to achieve acceleration.
We evaluate our method on the QwQ-32B model using the widely used benchmark AIME24 dataset, with a maximum output length set to 32k tokens. The results, illustrated in Figure 6, demonstrate a significant improvement in both generation speed and average acceptance tokens. Specifically, our method achieves a generation rate of 42.63 tokens/s, 2.25$\times$ higher than the strong Flash Attention baseline’s 18.92 tokens/s, and an average of 3.82 average acceptance tokens. Notably, QwQ-32B with LongSpec achieves even lower latency than the standard 7B model with Flash Attention, demonstrating that our method effectively accelerates the long reasoning model. These findings not only highlight the effectiveness of our method in the long reasoning task but also provide new insights into lossless inference acceleration for the o1-like model. We believe that speculative decoding will play a crucial role in accelerating this type of model in the future.
@@ 4.5 Throughput
As illustrated in Figure 7, the throughput results of Vicuna-7B on the RepoBench-P dataset show that LongSpec consistently outperforms both Vanilla and MagicDec across all batch sizes. At a batch size of 8, LongSpec achieves a throughput of 561.32 tokens/s, approximately 1.8$\times$ higher than MagicDec (310.58 tokens/s) and nearly 2$\times$ higher than Vanilla (286.96 tokens/s). MagicDec, designed with throughput optimization in mind, surpasses Vanilla as the batch size increases, reflecting its targeted improvements. However, LongSpec still sustains its advantage, maintaining superior throughput across all tested batch sizes.
@@ 5 Conclusion
In this paper, we propose LongSpec, a novel framework designed to enhance lossless speculative decoding for long-context scenarios. Unlike previous speculative decoding methods that primarily focus on short-context settings, LongSpec directly addresses three key challenges: excessive memory overhead, inadequate training for large position indices, and inefficient tree attention computation. To mitigate memory constraints, we introduce an efficient draft model architecture that maintains a constant memory footprint by leveraging a combination of sliding window self-attention and cache-free cross-attention. To resolve the training limitations associated with short context data, we propose the Anchor-Offset Indices, ensuring that large positional indices are sufficiently trained even within short-sequence datasets. Finally, we introduce Hybrid Tree Attention, which efficiently integrates tree-based speculative decoding with Flash Attention. Extensive experiments demonstrate the effectiveness of LongSpec in long-context understanding tasks and real-world long reasoning tasks. Our findings highlight the importance of designing speculative decoding methods specifically tailored for long-context settings and highlight promising directions for future research in efficient large-scale language model inference.
Appendix A Related Work about Lossy Speculative Decoding
While original speculative decoding methods are mainly lossless, some recent works try to relax the constraints and explore lossy speculative decoding. For instance, BiLD (Kim et al., 2023) employs a small model for autoregressive text generation, with a larger model occasionally invoked non-autoregressively to refine inaccurate predictions, thereby achieving speedups with minimal quality degradation. Narasimhan et al. (Narasimhan et al., 2025) introduce speculative cascading, a method that integrates cascade-style deferral rules with speculative execution to yield better cost-quality trade-offs than either approach alone. Another approach, MTAD (Qin et al., 2025), uses a smaller auxiliary model to approximate the multi-token joint distribution of a larger model, enhancing both inference speed and output effectiveness by accepting a bounded error in this approximation. To address the rejection of high-quality but non-aligned draft tokens, Bachmann et al. (Bachmann et al., 2025) propose adapting the verification step by training a compact “judge” module to recognize valid continuations even without perfect target model alignment, significantly boosting acceptance rates and speed. RSD (Liao et al., 2025) incorporates a process reward model to evaluate intermediate decoding steps, dynamically deciding target model invocation and introducing a controlled bias towards high-reward outputs to optimize the cost-quality trade-off. RAPID (Chen et al., 2025a) employs a RAG drafter on shortened contexts as a drafter. TokenSwift (Wu et al., 2025) comprehensively uses LLMs with partial KV cache and $N$-gram tables to accelerate ultra-long sequence generation (up to 100k tokens) while reducing computation time from hours to minutes.
Appendix B The Intuition of Why KV Cache Can Help
The KV cache stores the contextual information the model accumulates while processing previous tokens. When predicting the next token, the target model relies on three components: the KV cache (contextual memory), input word embeddings, and model parameters.
In our method, the draft model already shares the input embeddings with the target model, so the primary differences in their predictions stem from the KV cache and internal parameters. By allowing the draft model to use the KV cache generated by the target model, we eliminate another source of variation. As a result, the only remaining difference between their predictions comes from the model parameters. This sharing aligns the draft model’s predictions more closely with those of the target model, as it removes discrepancies caused by differing contextual representations.
Appendix C Correctness for Attention Aggregation